. A fair die is rolled six times. The probability that the highest outcomeof the six throws is less than 4 is 1/6. l . A CRV has uniform distribution from 5 to 3. The probability that the CRV is positive (F)11p (T is 3/8. h. A jar contains 3 red marbles, 7 green marbles and 10 white marbles. If a marble is drawn from the įarat random, the probability that this marble is white-os ip The slope of the PDF for CRV couldat be negative

True or False With explanation please.
Answer #1

f:

Out of six outcomes of a single throw of a dice, 3 are less than 4. That is the probability of getting a outcome less than 4 in a single roll is

P(less than 4) = 3/6 = 1/2

The probability that all outcomes in 6 throw are less than 4, that is highest outcomes of 6 throws will be less than 4, is

left ( rac{1}{2} ight )^{6}=rac{1}{64}

Hence, it is false.

g:

The probability that the CRV is positive is

3 3-(-5) L8J08

Hence, it is true.

h:

Total number of marbles in the jar = 3+7+10 = 20

Out of 20 marbles, 10 are white so

P(white) = 10 /20 = 0.50

Hence, it is true.

i:

Slope of CDF cannot be negative.

False